OpenStudy (anonymous):

!"# Δ! → 0 Still need help! lim = sec (x + Δx) − sec (x) Δx--0 Δx

OpenStudy (anonymous):

you will have to retype this in the comments using the equation button I dont even know exactly what is going on

OpenStudy (anonymous):

are you saying\[\lim_{\Delta x \rightarrow 0}=\frac{ \sec(x+\Delta x)-\sec(x) }{ \Delta x }\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok so first multiply by the conjugate of \[\sec(x+\Delta x)+\sec(x)\] what do you get

OpenStudy (anonymous):

sec^2(x + delta x) - sec^2 x

OpenStudy (anonymous):

or wait sorry still in the mindset of this other problem I have to take another look at this

OpenStudy (anonymous):

change your sec to 1/cos and simplify that

OpenStudy (anonymous):

I changed them to 1/cos and got a common denominator to combine them \[cosx - \cos(x + Deltax) \over \cos(x + deltax)cosx\]

OpenStudy (anonymous):

I didn't think I could cancel the cos because they are being subtracted in the numerator. Did I miss something?

OpenStudy (anonymous):

I dont understand what you did when you changed to 1/cos because you would have\[\frac{ 1 }{ \cos(x+\Delta x)-\cos(x) }*\frac{ 1 }{ \Delta x }\]

OpenStudy (anonymous):

and sorry you were right about the cos

OpenStudy (anonymous):

@agent0smith we need your help

OpenStudy (anonymous):

yeah I'm getting stumped myself

OpenStudy (anonymous):

\[\frac{ 1 }{ cosx + \delta x } - \frac{ 1 }{ cosx }\times \frac{ 1 }{ Deltax }\] Then I got a common denominator between the two fractions.

OpenStudy (anonymous):

oh wait I see now

OpenStudy (anonymous):

since it is as dleta x approaches zero you can plug in zero for delta x now and simplify

OpenStudy (anonymous):

it should simplify the fraction more for you

OpenStudy (anonymous):

Will that give me zero in the denominator?

OpenStudy (anonymous):

plug it in\[\frac{ \cos(x)-\cos(x-0) }{ \cos(x+0)\cos(x) }\]

OpenStudy (anonymous):

so you now have \[\frac{ \cos(x)-\cos(x) }{ \cos ^{2}(x) }\]

OpenStudy (anonymous):

I still have the 1/delta x also from the original problem since it was all over delta x to begin with. If I replace delta x with zero, does that give me zero in the denominator?

OpenStudy (anonymous):

show me where it would be in the equation I'm really sorry but I am cooking dinner at the same time so you will have to do most of the heavy lifting

OpenStudy (anonymous):

Bless your heart for doing calculus AND dinner! The original problem was completely over delta x. We inverted and multiplied leaving a delta x in the denominator. I was hoping that somehow the delta x's would cancel out, but I have tried every identity and trick I know and cannot get rid of them.

OpenStudy (agent0smith):

I'm guessing so far you've changed the secants into 1/cosine, and then used the angle subtraction formulas... i'll work it out myself though

OpenStudy (anonymous):

thanks agent0smith we need your help we've kind of been all over the place

OpenStudy (anonymous):

I used the angle addition of Cos(A +B) = CosACosB -SinASinB --then I kind of got stuck.

OpenStudy (agent0smith):

After some simplifying it looks like \[\Large \frac{ \cos x - \cos x \cos \Delta x+ \sin x \sin \Delta x }{ \cos x( \cos x \cos \Delta x - \sin x \sin \Delta x) } * \frac{ 1 }{\Delta x }\] The result should eventually be secxtanx, since we're finding the derivative of secx...

OpenStudy (anonymous):

I got this far as well, but I must be missing something in the simplification of this. How did you get to secxtanx? Did you split it into two expressions and work them separately?

OpenStudy (agent0smith):

I didn't get to secxtanx. I said that the derivative of secx is secxtanx.

OpenStudy (anonymous):

OH..sorry, I misunderstood. What is the limit of this problem as delta x approaches zero?

OpenStudy (agent0smith):

I haven't figured out how to show the work yet, but the original problem is basically finding the derivative of secx (you might be too early into calculus/precalc to know this yet), so your eventual solution will be secxtanx.

OpenStudy (anonymous):

Thank you so much for your help. I haven't had calculus for a lot of years and I am trying to look up what I do not remember! I am sure it will come back to me with some practice. If you think of a way to show how the problem simplifies--that would be great. Thanks again to both of you recon14193 and Agent0smith

OpenStudy (agent0smith):

I tried using wolfram alpha to simplify it, no help :/

OpenStudy (agent0smith):

Hmm, if i divide the numerator by cosx i'll get a tanx up top \[\Large \frac{ \cos x - \cos x \cos \Delta x+ \sin x \sin \Delta x }{ \cos x( \cos x \cos \Delta x - \sin x \sin \Delta x) } * \frac{ 1 }{\Delta x }\]

OpenStudy (agent0smith):

After dividing the numerator and denominator by cos x \[\Large \frac{1 - \cos \Delta x+ \tan x\sin \Delta x }{ \cos x \cos \Delta x - \sin x \sin \Delta x } * \frac{ 1 }{\Delta x }\]

OpenStudy (agent0smith):

Still don't see the Delta x disappearing anytime soon but i'll try to figure it out....

OpenStudy (agent0smith):

Maybe this will help http://www.youtube.com/watch?v=ClGVHlpx9_A

OpenStudy (anonymous):

now lets rewrite \[\lim_{\Delta x \rightarrow 0}\left( \frac{ 1-\cos \Delta x +\tan x \sin \Delta x}{ \cos x \cos \Delta x-\sin x \sin \Delta x }\frac{ 1 }{ \Delta x } \right)=\frac{ \lim_{\Delta x \rightarrow 0}\frac{ 1-\cos \Delta x+\tan x \sin \Delta x }{ \Delta x } }{ \lim_{\Delta x \rightarrow 0}\left( \cos x \cos \Delta x -\sin x \sin \Delta x\right) }\] now you can figure out the denominator by substituting 0 Delta x, cos(0)=1 and sin(0)=0, so denominator becomes 1/cos(x) or sec(x). now lets figure out top equation:\[\lim_{\Delta x \rightarrow 0}\frac{ 1-\cos \Delta x+\tan x \sin \Delta x }{ \Delta x }=\lim_{\Delta x \rightarrow 0}\frac{ 1-\cos \Delta x }{ \Delta x }+\tan x \lim_{\Delta x \rightarrow 0}\frac{ \sin \Delta x }{ \Delta x}=\tan x\] because \[\lim_{\Delta x \rightarrow 0}\frac{ 1-\cos \Delta x }{ \Delta x }=0\] and\[\lim_{\Delta x \rightarrow 0}\frac{ \sin \Delta x }{ \Delta x }=1\] so your limit becomes \[\sec x \tan x\]

OpenStudy (anonymous):

if you need help prooving those 2 limits i gave you answers, I'm here to help

OpenStudy (agent0smith):

Is there anyway to prove lim (1-cosx)/x = 0 w/o l'hopitals rule? That's what i was trying to avoid, any use of that (i feel like otherwise you can just use it right from the start). Good solution regardless.

OpenStudy (anonymous):

yes there is\[\lim_{\Delta x \rightarrow 0}\frac{ 1- \cos \Delta x }{ \Delta x }=\lim_{\Delta x \rightarrow 0}\left( \frac{ 1-\cos \Delta x }{ \Delta x }\frac{ 1+\cos \Delta x }{ 1+\cos \Delta x } \right)=\lim_{\Delta x \rightarrow 0}\frac{ 1-\cos^2 \Delta x }{ \Delta x \left( 1+\cos \Delta x \right) }=\lim_{\Delta x \rightarrow 0}\frac{ \sin^2 \Delta x }{ \Delta x \left( 1+\cos \Delta x \right) }=\lim_{\Delta x \rightarrow 0}\frac{ \sin \Delta x }{ \Delta x }\lim_{\Delta x \rightarrow 0}\frac{ \sin \Delta x }{ 1+\cos \Delta x }\]

OpenStudy (anonymous):

and you prove sin(Delta x)/(Delta x)=1 by squeezing theorem

OpenStudy (anonymous):

l'hopital's rool is easiest way out

OpenStudy (anonymous):

or just differentiating 1/cosx by chain rool